# ViTally Consistent: Scaling Biological Representation Learning for Cell Microscopy

### v1

@@ 5 Whole-genome benchmarking
Similarly, linear probing to select optimal ViT blocks led to significant improvements even when applied to frozen Dino-V2 based models pretrained on natural images. Dino-V2 ViT-G obtains a nearly 20% improvement CORUM recall (.44$\rightarrow$.53) by using the embeddings extracted at $b^{*}=16$ (chosen by linear probes) rather than the final embedding from $b=40$ (which performs worse than a random untrained ViT-S). Dino-V2 ViT-S also observes improvements by using $b^{*}=5$ rather than $b=12$ and outperforms Dino-V2 ViT-G in replicate consistency.
Figure 5: Replicate consistency whole-genome results between perturbations from cosine similarity distributions on RxRx3 post-TVN. Comparing baseline Dino-V2 ViT-G/14 at its typically used final block (left) versus MAE-G/8 at the best block found via linear probing (right).
@@ 6 Discussion and Conclusions
This work demonstrates that: (1) within the context of biological imaging, trimming many ViTs to an earlier block leads to stronger biological linearity and improved performance on downstream tasks in addition to cheaper inference costs (Figure 3); (2) linear probing performance on a subset of genetic perturbations correlates strongly with downstream performance on whole-genome benchmarks and can be used to optimize which block is selected for representing the whole-genome (Figure 4); (3) the most scaled model, MAE-G/8 , obtains the overall best performance across all benchmarks and linear probes, providing further evidence for the scaling hypothesis in biological image data (Table 2). This demonstrates that intentionally scaling training compute and parameters of SSL models for microscopy can benefit downstream biological relationship recall, whole-genome replicate consistency, and biological linear separability on smaller datasets (see § A.9 for scaling plots).
More broadly, this work proposes a reusable recipe for training and extracting optimal representations from fully self-supervised models trained on experimental data. The pattern we use can be applied to other domains that contain data from repeated experiments but without accurate ground truth labels. Specifically, we recommend: (1) curating the training set by identifying diverse sets of samples that are represented consistently, e.g., by using a pre-existing model to select such samples; (2) training a scaled transformer-based model using a self-supervised learning technique, such as masked autoencoding; and, (3) evaluating the performance of the trained transformer at every block to identify the optimal layer for representing the data.
@@ A.2 Training hyperparameters
Table 3: Training hyperparameters for the new models presented in this work. Each used a one-cycle cosine learning rate decay schedule with 10% warm-up using the Lion optimizer from Chen et al. (2023b) with betas (0.9, 0.95) and weight decay of 0.05, with additional ViT settings such as LayerScale as proposed by Dehghani et al. (2023). Note that MAE-G/8 had multiple restarts during training due to challenges associated with massive model training on large-scale shared distributed compute clusters.
@@
Table 3 provides the hyperparameters used for training the new vision transformers presented in this work. Each model was trained using a 75% mask ratio and the standard decoder architecture for MAEs (He et al., 2022). Each model was trained with the standard L2 MAE loss and the Fourier-space loss function implemented by Kraus et al. (2024) with a weight of $\alpha=0.01$. We note, however, that the details presented by Kraus et al. (2024) do not precisely correspond with the implementation provided in their Github repository; when reshaping the tokens to a shape compatible with the 2D Fourier transform, the permute operation resulted in adjacent pixels being from different channels of the input, resulting in the high frequency components of the loss being a function of the relationships between input channels. An initial investigation with a ViT-L/8 showed that changing the implementation to the one described in the paper did not dramatically change probing results. As such, we used the implementation as-is and leave additional analysis of loss function design for MAEs to future work.
@@ A.3 Perturbation Consistency
In order to assess the consistency of the induced morphology on the cells by the perturbations, we used a non-parametric perturbation consistency test similar to the one introduced in Celik et al. (2024). Let $x_{g,1},x_{g,2},\cdots,x_{g,n}$ be the embeddings for replicates of perturbation $x_{g}$ on experiment (batch) $e$. As the test statistic for perturbation consistency, $\bar{s}_{g}^{e}$ is defined as the mean of the cosine similarities across all pairs of replicates of $x_{g}$.
@@ A.4 Training linear probes
A.5 Anax classification for other cell lines/treatment conditions: ARPE19 and HUVEC with TNF-alpha background
@@
Figure 6: Layerwise validation set linear probe performance on Anax functional gene group classification beyond RxRx3: CRISPR knockouts in the ARPE-19 immortalized epithelial cell-line (left), and in HUVEC cells with a TNF-$\alpha$ background (right).
@@ A.6 Biological Relationship Recall
A valuable use of large-scale HCS experiments is to perform large-scale inference of biological relationships between genetic perturbations. We evaluate each model’s ability to recall known relationships by using the biological relationship recall benchmark described in Celik et al. (2024). First, we correct for batch effects using Typical Variation Normalization (TVN) (Ando et al., 2017b), and also correct for possible chromosome arm biases known to exist in CRISPR-Cas9 HCS data (Lazar et al., 2023). To infer biological relationships, we compute the aggregate embedding of each perturbation by taking the spherical mean over its replicate embeddings across experiments. We use the cosine similarity of a pair of perturbation representations as the relationship metric, setting the origin of the space to the mean of negative controls. We compare these similarities with the relationships found in the following public databases: CORUM (Giurgiu et al., 2019), hu.MAP (Drew et al., 2017), Reactome (Gillespie et al., 2021), and StringDB Szklarczyk et al. (2020) (with $>$95% combined score). Table 2 reports the recall of known relationships amongst the top and bottom 5% of all cosine similarities between CRISPR knockout representations in RxRx3 (Fay et al., 2023).
@@ A.7 Replicate Consistency
Since the pairs are randomly selected for $p^{e_{i}}_{0}$, the embeddings would be mostly orthogonal thus the distribution would be centered around 0, similar to what Figure 5 illustrates. Given that not all CRISPR knockouts would induce a morphological change in the cells, it’s plausible for distribution $q^{e_{i}}$ to exhibit a peak around 0. As the model approaches the precision of an oracle, we would anticipate the mass situated around this peak to shift towards higher cosine similarity values.
@@ A.8 Anax Group Prediction Details
A.9 Correlation between model scale and benchmark results
In Figure 7 we show the correlations between training FLOps (floating point operations) and downstream results. Over all benchmarks we observe a very strong consistent linear trend where scaling training FLOps improves overall pwerformance. This work provides the next log step in scale as we enter into the billion-parameter model regime with MAE-G/8. These results therefore provide additional evidence that the trend initially discovered by Kraus et al. (2023) between FLOps and relationship recall actually extends both to billion-parameter models and even moreso for other biologically meaningful benchmarks pertaining to linear probes on small experiments and to replicate consistency on the whole-genome.

### v2

@@ 5 Whole-genome benchmarking
Linear probing to select optimal ViT blocks leads to significant improvements even when applied to Dino-V2 and MAE models pretrained on natural images. Dino-V2 ViT-G is dramatically better in terms of recall and consistency with the early block embeddings at $b^{*}=16$ rather than the final embedding from $b=40$ (which performs worse than a random untrained ViT-S). Dino-V2 ViT-S also observes improvements by using $b^{*}=5$ rather than $b=12$ and outperforms Dino-V2 ViT-G in replicate consistency, while the trimmed Dino-V2 ViT-L obtains the best recall among the baselines. This finding is important for the scientific community to consider when applying ML foundation models to experimental data, as it is common to take embeddings from the final layer as a default strategy even when processing potentially out-of-distribution images (Lastufka et al., 2024), which, as we have shown, can significantly hinder results.
5.1 Performance on external data (JUMP-CP)
In order to validate that the MAEs generalize to entirely novel data, we evaluated a subset of models on completely external public data generated by different assays and from a variety of different labs as produced by the JUMP-CP consortium (Chandrasekaran et al., 2023). Table 2 presents these results using the relationship recall benchmarks of (Celik et al., 2024), noting that only a subset of 7,976 gene-knockouts are covered by this dataset. For post-processing embeddings, we use PCA with center-scaling for standard batch correction alignment. We observe that the MAEs perform better than the CellProfiler manual feature extraction baseline (§ A.11) (Carpenter et al., 2006; Kamentsky et al., 2011), and that the general trend is maintained with the trimmed MAE-G/8 obtaining the best recall overall. Notably, recall on JUMP-CP is considerably lower than on RxRx3 (Table 1) likely due to different assay protocols and more variance in the data.
6 RxRx3-core benchmarking
RxRx3-core [^2] https://huggingface.co/datasets/recursionpharma/rxrx3-core (Kraus et al., 2025) is a publicly available benchmarking dataset for assessing biological capabilities of computer vision models. RxRx3-core includes labeled images (compressed to JPEG-2000) of 735 genetic knockouts and 1,674 small-molecule perturbations across eight concentrations drawn from 222,601 wells ($512\times 512\times 6$ pixel center-crops) drawn from the larger RxRx3 dataset.
We evaluate a random choice baseline, CellProfiler, the CA-MAE-S/16 model, the MAE-L/8 model from previous work (Kraus et al., 2023), and the MAE-G/8. We evaluate both the trimmed and full-length version of the latter two models to determine the impact of our model-trimming strategy for the most performant models in this context.
We also evaluate a preliminary version of a SSL ViT-L/16 we trained with the Dino-V2 algorithm on the RxRx3 microscopy dataset; however, we found that it underperformed compared to the CellProfiler 4 baseline (§ A.11) (Kamentsky et al., 2011) and, unlike the MAEs we trained, significantly overfit during pretraining (§ A.9). As such, we did not pursue additional analysis on this model. We leave further development of Dino-V2 for microscopy data to future work, as we were unable to determine an effective recipe for applying Dino SSL pretraining on large-scale microscopy data.
To evaluate each model, we first inference all 222,601 $\times$ 4 crops and then average the 4 embeddings to the well-level. Then, to perform standard batch correction alignment, we use the “EMPTY”, unperturbed wells as our control population. We fit PCA on those control embeddings, use it transform the rest, and then fit a separate standard-scaler on each batch’s controls to transform the rest. This simplified alignment strategy empirically performed better than TVN on this dataset.
We present results for the benchmark measuring zero-shot prediction of compound-gene activity using cosine similarities between embeddings (Figure 5). This measures, for each compound, whether the cosine similarities from a model’s embeddings correctly rank the compound’s known target genes higher than a randomly sampled set of other genes from a ground truth dataset. Table 3 provides exact values along the max axis, which captures the strongest potential interaction regardless of concentration. The relative ranking of model performance, holds as expected from the results in § 4 and § 5, and trimming benefits both MAE-L/8 and MAE-G/8, with MAE-G/8 offering a 42% (3.77 $\rightarrow$ 5.38) relative improvement over the method from previous work in predicting compound-gene activity.
Table 3: Performance on public RxRx3-core compound-gene benchmark, measuring mean ($\pm$ STD over 100 random seeds) average precision in predicting compound activity against target genes with its z-score of improvement over the random baseline.
Model Avg. prec. $\uparrow$ Z-score $\uparrow$
Random baseline 0.222 $\pm$ .007 0.00
CellProfiler 0.274 $\pm$ .019 2.55
Dino-V2 ViT-L/16, RxRx3 0.258 $\pm$ .015 2.13
CA-MAE-S/16, RxRx3 0.273 $\pm$ .016 2.90
MAE-L/8, RPI-93M 0.290 $\pm$ .017 3.77
trimmed 0.299 $\pm$ .016 4.49
MAE-G/8, PP-16M 0.302 $\pm$ .015 4.79
trimmed 0.309 $\pm$ .015 5.38
Figure 5: Mean average precision performance on RxRx3-core public benchmark in predicting compound activity against annotated gene targets, across all compound concentrations with error bars for 100 runs of the benchmark with different random seeds (Table 3).
@@ 7 Discussion and Conclusions
This results in this work demonstrate that: (1) within the context of biological imaging, trimming many ViTs to an earlier block leads to stronger biological linearity and improved performance on downstream tasks in addition to cheaper inference costs (Figure 3); (2) linear probing performance on a subset of genetic perturbations correlates strongly with downstream performance on whole-genome benchmarks and can be used to optimize which block is selected for representing the whole-genome (Figure 4); (3) the most scaled model, MAE-G/8 trained on the specially curated dataset Phenoprints-16M for 500 epochs, obtains the overall best performance across all benchmarks and linear probes, providing further evidence for the scaling hypothesis in biological image data (Table 1, Table 2, Table 3, § A.10). We have found that intentionally scaling training compute and parameters of MAEs for microscopy on curated data can benefit a wide variety of biologically relevant tasks, even in comparison to a model trained on a 5×5$\times$5 × larger dataset. Indeed, even after training the MAE-G/8 on more than eight billion crops from a 16-million-image dataset, the validation reconstruction loss continued to improve. This suggests that increasing the model’s parameters further on this data may continue to yield improvements.
More broadly, this work proposes a reusable recipe for training and extracting optimal representations from fully self-supervised models trained on experimental data. The pattern we use can be applied to other domains that contain data from repeated experiments but without accurate ground truth labels. Specifically, we recommend: (a) curating the training set by identifying diverse sets of samples that are represented consistently, e.g., by using a pre-existing model to select such samples; (b) training a scaled transformer-based model using a self-supervised learning technique, such as masked autoencoding; and, (c) inspecting the performance of the trained transformer, and all baselines, at every block to identify the optimal layer for representing the data.
Limitations and reproducibility
In this work, we evaluated publicly available SSL ViTs as baselines and trained new MAE models trained on a specially curated microscopy dataset. Our preliminary attempts to train ViTs with Dino on this microscopy data encountered suboptimal performance (§ 6). Consequently, we allocated our time and compute budget to investigate scaling MAE to ViT-G/8 on curated data. We recognize the potential for other SSL pretraining regimes and fine-tuning strategies (Singh et al., 2023; Lehner et al., 2024; Hondru et al., 2024; Khan & Fang, 2024; Alkin et al., 2024) oriented for microscopy data to lead to future improvements on these tasks. We publicly release the inference, reconstruction visualization and benchmarking code [^3] https://github.com/recursionpharma/maes_microscopy, along with the full weights for CA-MAE ViT-S/16444https://huggingface.co/recursionpharma/OpenPhenom.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning in its applications to the sciences. There are many potential societal consequences of our work, especially relating to the discovery of new biological relationships between genes, potential drug treatments for diseases, and overall accelerating the process of drug development. At the same time, the predictions of these models are not guaranteed to be correct, which is why utmost care must be taken to validate safety and efficacy in orthogonal assays and early-stage clinical trials.
@@ A.2 Training hyperparameters
Model Name Parameters Blocks Model Dim Pretraining Data
Baselines
Untrained ViT-S/16 25M 12 384 N/A
Dino-V2 ViT-S/14 25M 12 384 Natural images
Dino-V2 ViT-L/14 307M 24 1024 Natural images
Dino-V2 ViT-G/14 1,100M 40 1536 Natural images
ViT-L/16 MAE 307M 24 1024 Imagenet-21k
MAEs for microscopy
CA-MAE-S/16 25M 12 384 RxRx3
MAE-L/8 307M 24 1024 RPI-93M
MAE-L/8 307M 24 1024 Phenoprints-16M
MAE-G/8 1,860M 48 1664 Phenoprints-16M
Table 5: Training hyperparameters for the new models presented in this work. Each used a one-cycle cosine learning rate decay schedule with 10% warm-up using the Lion optimizer from (Chen et al., 2023) with betas (0.9, 0.95) and weight decay of 0.05, with additional ViT settings such as LayerScale as proposed by (Dehghani et al., 2023). Note that MAE-G/8 had multiple restarts during training due to challenges associated with massive model training on large-scale shared distributed compute clusters.
@@
Table 5 provides the hyperparameters used for training the new vision transformers presented in this work. Each model was trained using a 75% mask ratio and the standard decoder architecture for MAEs (He et al., 2022). Each model was trained with the standard L2 MAE loss and the Fourier-space loss function implemented by (Kraus et al., 2024) with a weight of $\alpha=0.01$. We note, however, that the details presented by (Kraus et al., 2024) do not precisely correspond with the implementation provided in their Github repository; when reshaping the tokens to a shape compatible with the 2D Fourier transform, the permute operation resulted in adjacent pixels being from different channels of the input, resulting in the high frequency components of the loss being a function of the relationships between input channels. An initial investigation with a ViT-L/8 showed that changing the implementation to the one described in the paper did not dramatically change probing results. As such, we used the implementation as-is and leave additional analysis of loss function design for MAEs to future work.
@@ A.3 Perturbation Consistency
In order to assess the consistency of the induced morphology on the cells by the perturbations, we used a non-parametric perturbation consistency test similar to the one introduced in (Celik et al., 2024). Let $x_{g,1},x_{g,2},\cdots,x_{g,n}$ be the embeddings for replicates of perturbation $x_{g}$ on experiment (batch) $e$. As the test statistic for perturbation consistency, $\bar{s}_{g}^{e}$ is defined as the mean of the cosine similarities across all pairs of replicates of $x_{g}$.
@@ A.4 Training linear probes
[…]
@@ A.6 Biological Relationship Recall
A valuable use of large-scale HCS experiments is to perform large-scale inference of biological relationships between genetic perturbations. We evaluate each model’s ability to recall known relationships by using the biological relationship recall benchmark described in Celik et al. (2024). First, we correct for batch effects using Typical Variation Normalization (TVN) (Ando et al., 2017), and also correct for possible chromosome arm biases known to exist in CRISPR-Cas9 HCS data (Lazar et al., 2024). To infer biological relationships, we compute the aggregate embedding of each perturbation by taking the spherical mean over its replicate embeddings across experiments. We use the cosine similarity of a pair of perturbation representations as the relationship metric, setting the origin of the space to the mean of negative controls. We compare these similarities with the relationships found in the following public databases: CORUM (Giurgiu et al., 2019), hu.MAP (Drew et al., 2017), Reactome (Gillespie et al., 2021), and StringDB (Szklarczyk et al., 2020) (with $>$95% combined score). Table 1 reports the recall of known relationships amongst the top and bottom 5% of all cosine similarities between CRISPR knockout representations in RxRx3 (Fay et al., 2023).
@@ A.7 Replicate Consistency
Since the pairs are randomly selected for $p^{e_{i}}_{0}$, the embeddings would be mostly orthogonal thus the distribution would be centered around 0.Given that not all CRISPR knockouts would induce a morphological change in the cells, it’s plausible for distribution $q^{e_{i}}$ to exhibit a peak around 0. As the model approaches the precision of an oracle, we would anticipate the mass situated around this peak to shift towards higher cosine similarity values.
@@ A.8 Anax Group Prediction Details
Figure 7: Loss curve when training Dino-v2 ViT-L/16 on RxRx3.
A.9 Dino-V2 pretraining on microscopy data
We attempted to train two Dino-v2 models on microscopy data. One ViT-L/16 from scratch on RxRx3, and another attempt of fine-tuning the MAE-L/8 on RPI-93M with the Dino-v2 losses. They had the following hyperparameter settings which were tuned on another dataset: output-dim 65536, 2 global crops, 4 local crops, dino loss weight 1.0, koleo loss weight 0.1, ibot loss weight 1.0, Lion optimizer with max learning rate 1e-5, weight decay 0.05, betas 0.9 0.95, and cosine annealing. In both cases, we observed significant over-fitting of the loss from the start (Figure 7).
Table 6: RxRx3-core benchmarks for our initial attempts to train Dino-V2 models on microscopy data. The latter was finetuned from the MAE-L/8 trained on RPI-93M. Results compare to Table 3.
DinoV2 model Avg. Prec. Z-score
ViT-L/16 RxRx3 0.258 $\pm$ .015 2.13
ViT-L/8 (ft.) RPI-93M 0.255 $\pm$ .018 1.76
In Table 6 we show that both models fail to improve on the RxRx3-core benchmark metrics (i.e., z-scores over the random baseline) versus the CA-MAE-S/16 RxRx3 model which had Z-score of 2.90 on average precision for predicting compound activity. We have not found an effective recipe for training Dino on microscopy data. However, we note that theoretical evidence exists arguing that MAE learning is in some ways equivalent to contrastive learning (Hondru et al., 2024), so even if an appropriate Dino recipe is found it would remain to be seen if it differs substantially from MAEs for microscopy given the same training compute. As described in the Limitations section, we expect that future work would have to dedicate significant training ablations and creativity to determine the best possible training recipe for training Dino on microscopy data.
A.10 Correlation between model scale and benchmark results
In Figure 8 we show the correlations between training FLOps (floating point operations) and downstream results. Over all benchmarks we observe a very strong consistent linear trend where scaling training FLOps improves overall pwerformance. This work provides the next log step in scale as we enter into the billion-parameter model regime with MAE-G/8. These results therefore provide additional evidence that the trend initially discovered by (Kraus et al., 2023) between FLOps and relationship recall actually extends both to billion-parameter models and even moreso for other biologically meaningful benchmarks pertaining to linear probes on small experiments and to replicate consistency on the whole-genome.
A.11 CellProfiler features
CellProfiler bioimage analysis software (Carpenter et al., 2006; McQuin et al., 2018) was used to compute features using classical segmentation and feature extraction algorithms. Benchmarking results using CellProfiler features are reported for JUMP-CP (§ 5.1) and RxRx3-core (§ 6). JUMP-CP CellProfiler features were downloaded from https://cellpainting-gallery.s3.amazonaws.com/index.html#cpg0016-jump. CellProfiler features for RxRx3-core were computed using version 2.2.0 (Kamentsky et al., 2011). Single cells were segmented after applying illumination correction and shape, intensity, and texture features were extracted from each cell, resulting in 952 dimensional profiles. For RxRx3-core, we kept only cells in the center 512x512 crop of original 2048x2048 images and then mean aggregated them. 4,749 wells of the 222,601 RxRx3-core wells were missing CellProfiler features, so these were held out of the benchmark computations for all models on RxRx3-core.